A nation's real GDP increased from $17.8 trillion to $18.2 trillion in one year. In that same year, the nation's population increased from 326 million to 331 million. 


a. Calculate the nation's real GDP per capita in year 1. 

b. Calculate the nation's real GDP per capita in year 2. 

c. Use the Rule of 70 to determine how many years it would take this nations Real GDP per capita to double. 
Answer #1

Real GDP per capita is the ratio of real GDP to its population

Real GDP per capita in year 1 = 17. 8 trillion / 326 million

= 54601.23 dollars

Real GDP per capita in year 2 = 18.2 trillion / 331 million

= 54984.89 dollars

Growth rate of real GDP per capita = (54984.89 - 54601.23)*100/54601.23 = 0.7%

According to the rule of 70 it will take 70/0.7 = 100 years for the real GDP per capita to get doubled
Answer #2

SOLUTION :


a.


GDP per capita in year 1 

= 17.8*10^6 million $ / 326 million 

=  54601.23  ($) (ANSWER).


b. 


GDP per capita in year 2 

= 18.2*10^6 million $ / 321 million 

=  54984.89 ($) (ANSWER).


c.


GDP per capita growth rate = (54984 .89 / 54601.23) - 1 = 0.007027 = 0.7027%


No. of years required to double the GDP per capita as per rule of 70

= 70 / 0.7027

= 99.62 = 100  years approx. (ANSWER).



answered by: Tulsiram Garg